(A) Solve the following buffer problem twice: using the R-ICE table method using the Henderson-Hasselbalch equation...

  1. (A) Solve the following buffer problem twice:

    1. using the R-ICE table method

    2. using the Henderson-Hasselbalch equation

    Calculate the pH of a solution made by dissolving 5.15 g of hypochlorous acid, HOCl, (MM = 51.46 g/mol) and 15.09 g in sodium hyochlorite, NaOCl (MM = 75.46 g/mol) in 1.00 L of water solution. pKa = 7.52.
Answer #1

Ans :- pH = 7.82

Explanation :-

Number of moles = Given mass in g / Gram molar mass

So,

Number of moles of HOCl = Given mass in g / Gram molar mass of HOCl

= 5.15 g / 51.46 g/mol

= 0.100 mol

Also, Molarity = Number of moles / Volume of solution in L

So,

Molarity of HOCl = [HOCl]= 0.100 mol / 1.00 L

= 0.100 M

Similarly,

Number of moles of NaOCl = Given mass in g / Gram molar mass of NaOCl

= 15.09 g / 75.46 g/mol

= 0.200 mol

Also, Molarity = Number of moles / Volume of solution in L

So,

Molarity of NaOCl = [NaOCl]= 0.200 mol / 1.00 L

= 0.200 M

By using Henderson-Hasselbalch equation :

pH = pKa + log [Sal]/[Acid]

pH = 7.52 + log [NaOCl] / [HOCl]

pH = 7.52 + log 0.200 M / 0.100 M

pH = 7.52 + log 2.0

pH = 7.52 + 0.3010

pH = 7.82

Hence, pH of the buffer solution = 7.82